# DRIFT: Deep Reinforcement Learning for Intelligent Floating Platforms Trajectories

### v1

@@ Abstract
This investigation introduces a novel deep reinforcement learning-based suite to control floating platforms in both simulated and real-world environments. Floating platforms serve as versatile test-beds to emulate microgravity environments on Earth. Our approach addresses the system and environmental uncertainties in controlling such platforms by training policies capable of precise maneuvers amid dynamic and unpredictable conditions. Leveraging state-of-the-art deep reinforcement learning techniques, our suite achieves robustness, adaptability, and good transferability from simulation to reality. Our Deep Reinforcement Learning (DRL) framework provides advantages such as fast training times, large-scale testing capabilities, rich visualization options, and ROS bindings for integration with real-world robotic systems. Beyond policy development, our suite provides a comprehensive platform for researchers, offering open-access at https://github.com/elharirymatteo/RANS/tree/ICRA24.
@@ I INTRODUCTION
Across the globe, there has been exponential growth in the adoption of small satellites, including cubesats [1, 2]. Thanks to their cost-effectiveness, they are now extensively used for both commercial and scientific purposes. Consequently, several countries and their space administrations have actively invested in advancing small satellite technology over the past few decades. The surge in space missions is creating a growing demand to test and validate the flight software and hardware on the ground prior to employing them in space. These experiments aim to enhance the historically low success rates of missions in space [3].
@@
Improving the reliability and autonomy of the motion systems plays a key role in boosting mission success. Currently, the primary approach to enhancing autonomous navigation and control of such systems involves conducting performance tests. These tests help understand essential parameters and their relationships within the control scheme involving different sensors and actuators. This knowledge is pivotal for the design and operation of these systems, contributing significantly to mission success rates. To emulate free-floating and free-flying satellite motion, a common solution is to use floating platforms: a rigid structure floating on top of an extremely flat and smooth surface using air bearings [4]. This allows 2D motions with very low friction, effectively replicating space-like conditions in a plane.
The traditional methods of aerospace vehicle trajectory planning primarily rely on optimal control techniques, which involve the derivation of open-loop control solutions based on system models and predefined objectives. This approach is intricate and often necessitates specialized optimization software to find flight paths adhering to various constraints. However, as aerospace vehicles regularly encounter state disturbances and uncertainties, there is a growing need for more robust and adaptable control strategies.
@@
In response to this challenge, this paper introduces a novel approach that harnesses the power of Deep Reinforcement Learning (DRL) to control a floating platform (FP) within a 2D environment. The utilization of DRL offers an alternative to the conventional deterministic and expert-driven control methods prevalent in aerospace trajectory planning. DRL involves a goal-oriented agent that interacts with its environment, learning control policy approximations. Importantly, this learning process enables the agent to handle stochastic events in the environments by exploring the state-control space using reward signals.
@@
Our main contributions lie in three key areas. First, we enhance RANS [5], a simulator previously developed by our team, to accommodate more complex tasks and a diverse range of environment randomization profiles. Second, we demonstrate the high-performance capabilities of the Proximal Policy Optimization (PPO) [6] algorithm in both simulation and experimental scenarios. We evaluate its effectiveness in accomplishing three distinct tasks: navigating to a specific position, achieving a desired pose, and tracking a target velocity. Finally, we conduct a comprehensive comparison between the PPO-based approach and traditional optimal control algorithms, such as the Linear Quadratic Regulator (LQR), showcasing the benefits and drabacks of the two methods under different environmental conditions.
Throughout this paper, we will refer to our robotic system, using Floating Platform (FP) and air-bearing platform interchangeably.
@@ II-A Floating Platforms control
Floating platforms are systems with air bearings attached to their lower surface. These bearings release pressurized air and create a thin film to levitate the platform, thereby counterbalancing its weight to produce a microgravity effect (in-plane components of gravity on the main body are negligible), thus emulating the drag-free and weightless environment of orbital spaceflight.
@@
When focusing solely on the maneuvers of free-floating platforms, noteworthy developments have emerged. For instance, an innovative 3D-printed platform named “Slider” has been introduced [7]. Slider, equipped with eight thrusters, which can be precisely controlled through either motion in one of the 4 cardinal directions or 2 rotations. Furthermore, [8] presents an extensive characterization of air bearings platforms while also introducing a vision-based navigation system.
@@ II-B Deep Reinforcement Learning for thrust-based control
Over the past few years, there has been a flourishing utilization of ML and RL applied to space Guidance Navigation and Control (GNC) problems. These applications encompass a wide range of tasks, including planetary landing [16], path planning for lunar or asteroid hopping rovers [17, 18], spacecraft orbit control within unknown gravitational fields [19], and spacecraft map generation during orbits around small celestial bodies [20]. However, it is worth noting that most research employing deep reinforcement learning for aerospace control tasks shows simulation results.
To the best of our knowledge, only one prior work, namely [21], used reinforcement learning to guide both simulated and real-world floating platforms. In this paper, the authors combine DRL as a guidance policy whose trajectories are fed to a conventional controller to track. This work provides guidance techniques that successfully output velocity signals for the simulation and the experimental facility, achieving comparable performance to that observed during training.
Our work distinguishes itself from [21] in several key aspects. Firstly, we introduce a DRL agent capable of directly controlling the output thrust of the floating platform, eliminating the need for a separate trajectory tracker scheme. Secondly, our primary focus lies in delivering a comprehensive framework for training, assessing, and benchmarking DRL agents and optimal control methods across a spectrum of environmental conditions. These conditions represent a significant source of complexity when deploying AI agents on real systems, especially within the demanding space environment.
@@ III-A Problem Formulation
We formulate the problem of controlling a FP’s maneuvers as a sequential decision-making task, wherein actions are selected to guide the system from an initial state to a target pose within a 2D space. The system considered in our experiments is a lightweight FP [22] developed at the University of Luxembourg. It is actuated by 8 “on-off thrusters” with a maximum thrust of 1N. The thrusters all share the same pressure line, such that when more than 1 thruster is fired, they all share equally the maximum available thrust. We consider a 10-dimensional state space defined as shown in Equation 1, where, at each time step $t$, $\theta$ represents the heading orientation, $v_{x}$ and $v_{y}$ are the linear velocities, $\omega_{z}$ is the angular velocity, tf denotes the task flag (integer to indicate which is the addressed task), and $\text{td}_{1-4}\in R$ include additional task-related data (such as distances to target position and target heading).
$s_{t}=(\cos(\theta),\sin(\theta),v_{x},v_{y},\omega_{z},\text{tf},\text{td}_{1},\text{td}_{2},\text{td}_{3},\text{td}_{4})$ (1)
The task-specific observations ($to$) are shown in Table I, where $\Delta$ denotes the distance between each variable (position, velocity or angle) and their target value. The observation space is structured as such so that this work can easily be extended to support multitask-capable policies. For the control of the platform, our agents use an 8-dimensional action space that corresponds to a binary activation of the available thrusters.
TABLE I: State task-specific observations
Task tf $\text{to}_{1}$ $\text{to}_{2}$ $\text{to}_{3}$ $\text{to}_{4}$
Go to position 0 $\Delta x$ $\Delta y$
@@
To train our agents, we used the reward functions defined in (234). These rewards are shaped using an exponential term, as we empirically found that models trained with them converged faster while being more accurate.
  • Go to position:
    $R=\exp\left(-\frac{e_{p}}{0.25}\right)-p$ (2)
  • Go to pose:
    $R=\exp\left(-\frac{e_{p}}{0.25}\right)\cdot S_{p}+\exp\left(-\frac{e_{o}}{0.25}\right)\cdot S_{o}-p$ (3)
  • Track velocity:
    $R=\exp\left(-\frac{e_{v}}{0.25}\right)-p$ (4)
Here, the errors are considered as the absolute distance to the relative target. $e_{v}$ is the velocity error, $e_{p}$ and $e_{o}$ are the position and orientation errors, and $S_{p}$ and $S_{o}$ their respective scaling coefficients. In our experiments, we set both $S_{p}$ and $S_{o}$ equal to 1. The quantity $p$ stands for the possible three additional penalties, which can be added to penalize energy consumption (high number of thrusters activation), high linear, and high angular velocities. We experimented with different penalty shapes and settled on using a linear penalty on the energy consumption as defined in (5).
$p=0.3\sum_{i=1}^{8}T_{i}$ (5)
@@ III-B Simulation
Building upon our prior simulator RANS [5], we introduce enhancements to enable the platform to perform more complex tasks. RANS leverage Nvidia’s IsaacSim, specifically relying on OmniIsaacGym [23], a versatile simulator, capable of concurrently running thousands of environments. In the original RANS framework, only nominal system and environmental conditions were present, limiting the agents to learn on ideal conditions, which are usually far from those encountered when using the real FP systems. To mitigate this gap, we introduce RANS v2.0 which includes the following extensions: possibility to easily fine-tune the control policies through parameterized rewards and penalties, a disturbance generator that allows the injection of action and state noises, of a force and torque disturbance acting on the system, and a random thruster deactivation. RANS v2.0, with only approximately 10 minutes required to train a model on an RTX 4090, achieves a throughput of more than 40,000 steps per second with all disturbances enabled, which is very close to its previous version. Furthermore, it enables large-scale testing by swiftly evaluating thousands of initial conditions in seconds, provides extensive built-in examples and task sets. It offers rich visualization options, including metric tracking during training through the WandB API [24], and comprehensive evaluation metrics presented through tables and plots. The library uses the OpenAI Gym [25] format to define the RL loop, including the standard normalization of the observation space. Additionally, the integration of a ROS interface enhances the versatility of our framework, allowing easy integration and deployment of the control policies within real-world robotic systems.
@@ III-C Training Procedure
We reworked the PPO implementation from the RL Games library [26] as the foundation of our training procedure. This implementation utilizes GPU acceleration to vectorize observations and actions, facilitating efficient parallelization within the simulator by having both the physics simulation and the policy training residing on GPU. Our agents are designed as actor-critic networks with two hidden layers, each consisting of 128 units. They undergo training within their respective environments for a total of 2000 epochs or approximately 130M steps. For more details about the network or the PPO configuration, we invite the curious reader to refer to the training configuration files available along with the code release.
III-D Benchmark comparison
@@
In this paper, we aim to provide a benchmark comparison between deep reinforcement learning and optimal control approaches, LQR in particular, for addressing the control problem of the floating platform in various scenarios. Our objective is not to establish the superiority of one method over the other, but rather to gain insights into the strengths and weaknesses of each approach under different environmental conditions and task requirements. To facilitate this comparison, we replicate the environment we had inside Isaac in MuJoCo [27]. We found MuJuCo more practical to evaluate the performance of traditional control algorithms, as it can be implemented easily without any parallelization or PyTorch dependencies. It also enabled us to test the policies outside the Omniverse environment and prepare for sim-to-real transfer. Our evaluation aims to provide insights into the relative strengths and weaknesses of RL and traditional control methods in our problem.
III-E Optimal Controller
An infinite horizon discrete-time LQR controller [28] is used as a preliminary comparison with the DRL algorithm to control the FP. The position, linear velocities, orientation quaternions, and angular velocities in the 2D plane are considered as state variables of the system. The corresponding state matrices are calculated using finite differencing in MuJoCo so that they correspond to the control outputs from the eight thrusters. To account for the disturbances endured by the FP, the system matrices are updated at regular intervals.
III-F Laboratory Experiment Setup

### v2

@@ Abstract
This investigation introduces a novel deep reinforcement learning-based suite to control floating platforms in both simulated and real-world environments. Floating platforms serve as versatile test-beds to emulate microgravity environments on Earth, useful to test autonomous navigation systems for space applications. Our approach addresses the system and environmental uncertainties in controlling such platforms by training policies capable of precise maneuvers amid dynamic and unpredictable conditions. Leveraging Deep Reinforcement Learning (DRL) techniques, our suite achieves robustness, adaptability, and good transferability from simulation to reality. Our deep reinforcement learning framework provides advantages such as fast training times, large-scale testing capabilities, rich visualization options, and ROS bindings for integration with real-world robotic systems. Being open access, our suite serves as a comprehensive platform for practitioners who want to replicate similar research in their own simulated environments and labs.
@@ I INTRODUCTION
Across the globe, there has been an exponential growth in the adoption of small satellites, including cubesats [1, 2]. Thanks to their cost-effectiveness, they are now extensively used for both commercial and scientific purposes. Consequently, several countries and their space administrations have actively invested in advancing small satellite technology over the last few decades. The surge in space missions is creating a growing demand to test and validate the flight software and hardware on the ground prior to employing them in space. These experiments aim to enhance the historically low success rates of missions in space [3].
@@
Improving the reliability and autonomy of the motion systems plays a key role in boosting mission success. Currently, the primary approach to enhancing autonomous navigation and control of such systems involves conducting performance tests. These tests help understand essential parameters and their relationships within the control scheme involving different sensors and actuators. This knowledge is pivotal to the design and operation of these systems, contributing significantly to mission success rates [1]. To emulate free-floating and free-flying satellite motion, a common solution is to use floating platforms: a rigid structure floating on top of an extremely flat and smooth surface using air bearings [4]. This allows 2D motions with very low friction, effectively replicating space-like conditions in a plane.
@@
The traditional methods of aerospace vehicle trajectory planning primarily rely on optimal control techniques [3], which involve the derivation of open-loop control solutions based on system models and predefined objectives. These approaches are intricate and often necessitate specialized optimization software to find flight paths adhering to various constraints. However, as aerospace vehicles regularly encounter state disturbances and uncertainties, there is a need for more robust and adaptable control strategies. In response to this challenge, this paper introduces a novel approach that harnesses the power of Deep Reinforcement Learning (DRL) to control a floating platform (FP) within a 2D environment. The utilization of DRL offers an alternative to the conventional deterministic and expert-driven control methods prevalent in aerospace trajectory planning. DRL involves a goal-oriented agent that interacts with its environment, learning control policy approximations. Importantly, this learning process enables the agent to handle stochastic events in the environments by exploring the state-control space using reward signals.
Figure 1: Floating platform in ZeroG Laboratory.
@@
Our main contributions lie in three key areas. First, we enhance RANS [5], a simulator previously developed by our team, to accommodate more complex tasks and a diverse range of environment randomization profiles. Second, we demonstrate the high-performance capabilities of the Proximal Policy Optimization (PPO) [6] algorithm in both simulated and real scenarios. We evaluate its effectiveness by completing two distinct tasks: navigating to a specific position and orientation, and tracking a target velocity. Finally, we conduct a comprehensive comparison between the PPO-based approach and traditional optimal control algorithms, such as the Linear Quadratic Regulator (LQR) [7], showcasing the benefits and drabacks of the two methods under different environmental conditions.
@@ II-A Floating Platforms control
Floating platforms are systems with air bearings attached to their lower surface. These bearings release pressurized air creating a thin film to levitate the platform; thereby counterbalancing its weight to produce a microgravity effect (in-plane components of gravity on the main body are negligible), thus emulating the friction-less and weightless environment of orbital spaceflight.
@@
When focusing solely on the maneuvers of free-floating platforms, noteworthy developments have emerged. For instance, an innovative 3D-printed platform named “Slider” has been introduced [8]. Slider, equipped with eight thrusters, can be precisely controlled through either motion in one of the 4 cardinal directions or 2 rotations. Furthermore, [9] presents an extensive characterization of air bearings platforms while introducing a vision-based navigation system that takes into account the vibrations caused by the thrusters.
@@ II-B Deep Reinforcement Learning for thrust-based control
Over the past few years, there has been a flourishing utilization of ML and RL applied to space Guidance Navigation and Control (GNC) problems. These applications encompass a wide range of tasks, including: planetary landing [17]; path planning for lunar or asteroid hopping rovers [18, 19]; spacecraft orbit control within unknown gravitational fields [20]; and spacecraft map generation during orbits around small celestial bodies [21]. However, it is worth noting that most research employing deep reinforcement learning for aerospace control tasks shows numerical simulation results only.
To the best of our knowledge, only two prior works [22, 23], used reinforcement learning to guide both simulated and real-world floating platforms. In [22], the authors combine DRL as a guidance policy whose trajectories are fed to a conventional controller to track. This work provides guidance techniques that successfully output velocity signals for the simulation and the experimental facility, achieving comparable performance to that observed during training. In [23] the main focus is on reaching a specific target while avoiding static or dynamic obstacles.
Our work distinguishes itself from [22] and [23] in several key aspects. Firstly, we introduce a DRL agent capable of directly controlling the output thrust of the floating platform, eliminating the need for a separate trajectory tracker scheme. Secondly, our primary focus lies in delivering a comprehensive framework for training, assessing, and bench-marking DRL agents and optimal control methods across a spectrum of environmental conditions. These conditions represent a significant source of complexity when deploying AI agents on real systems, especially within the demanding space environment.
@@ III-A Problem Formulation
Figure 2: Floating platform and target in global reference frame.
In this paper, we approach the task of guiding a FP’s maneuvers as a sequential decision-making problem. To facilitate our investigation and demonstrate the practical applicability of our proposed techniques—from sim to real-world scenarios, we simplify the complex orbital dynamics into a two-dimensional kinematic model. As illustrated in Figure 2, we use a global reference frame (denoted $W$). This allows for consistent and absolute measurements of the position and heading errors. The framework also allows for the use of local coordinates whenever considered convenient.
Within this framework the control policy must learn the optimal sequence of actions by observing state transitions, thereby minimizing the task-specific error. We define the different tasks as: (i) Go to pose, starting from a random initial position in the plane, reach the given pose (position and orientation $\theta$); (ii) Track velocity, track the given velocity vector, which can in turn be used to follow a trajectory.
For both tasks the control policy is required to minimize the error metrics derived from the current state observations of the floating platform and the target. Regarding the “go to pose” task, the positional error is defined as the Euclidean distance between the FP’s current position, $p_{fp}=(x_{fp},y_{fp})$, and the target position, $p_{t}=(x_{t},y_{t})$, Eq. (1), while the heading error is calculated based on the difference between the platform’s current orientation $\theta_{fp}$ and the target heading $\theta_{t}$, Eq. (2):
$e_{p}$ $=\|\mathbf{p}_{\text{t}}-\mathbf{p}_{\text{fp}}\|_{\scriptscriptstyle 2}$ (1)
$e_{\theta}$ $=\arctan 2\left(\sin(\theta_{\text{t}}-\theta_{\text{fp}}),\cos(\theta_{\text{t}}-\theta_{\text{fp}})\right)$ (2)
For the “track velocity” task the angular and linear velocity errors ($\mathbf{e}_{v},\mathbf{e}_{\omega}$) are determined by subtracting the FP’s current velocities ($\mathbf{v}_{fp}$) from the target velocities ($\mathbf{v}_{t}$), Eq. (3) and (4).
$\mathbf{e}_{v}$ $=\mathbf{v}_{\text{t}}-\mathbf{v}_{\text{fp}}$ (3)
$\mathbf{e}_{\omega}$ $=\boldsymbol{\omega}_{\text{t}}-\boldsymbol{\omega}_{\text{fp}}$ (4)
In our study, we focus on a lightweight floating platform (FP) developed at the University of Luxembourg [24]. The system is defined by a 10-dimensional state space, Eq. 5. At each discrete time step $t$, the state variables include the FP’s heading ($\theta$), its linear velocities ($v_{x}$ and $v_{y}$), angular velocity ($\omega_{z}$), a task flag (f) indicating the current task, and four additional variables ($\text{d}_{1-4}$) representing task-specific data such as distances to the target position and heading:
$s_{t}=(\cos(\theta),\sin(\theta),v_{x},v_{y},\omega_{z},\text{f},\text{d}_{1},\text{d}_{2},\text{d}_{3},\text{d}_{4})^{\top}.$ (5)
Task-specific data, written $\text{d}_{1-4}$, is detailed in Table I, where $\Delta$ denotes the vector norm distance between the variables (such as position, velocity, or angle) and their respective target values. This configuration of the observation space is intentionally designed to facilitate the future extension of this work to learn policies capable of handling multiple tasks simultaneously.
For the control of the platform, our agents use an 8-dimensional action space that corresponds to a binary activation of 8 “on-off thrusters”. These share the same pressure line, such that, at every step of the control loop, the maximum force generated by each thruster is $\frac{1}{n}$ N where $n$ is the number of active thrusters. Simply put, if only one thruster is turned on, it will output 1 Newton, if 2 thrusters are activated they generate 0.5 N each, etc.
TABLE I: State task-specific data.
Task f $\text{d}_{1}$ $\text{d}_{2}$ $\text{d}_{3}$ $\text{d}_{4}$
@@
To guide the optimization process for the control policies, an exponential reward structure was adopted, as after empirical evaluation it was found to yield faster and more accurate convergence. In particular, Eq. (6) for the “go to pose” task and Eq. (7) for the “track velocity” task were used:
$R_{po}$ $=\exp\left(-\frac{e_{p}}{0.25}\right)\cdot S_{p}+\exp\left(-\frac{e_{\theta}}{0.25}\right)\cdot S_{\theta}-p$ (6)
$R_{v}$ $=\exp\left(-\frac{e_{v}}{0.25}\right)\cdot S_{p}+\exp\left(-\frac{e_{\omega}}{0.25}\right)\cdot S_{\theta}-p$ (7)
In this context, errors are quantified as the norm distance from the specified targets, with $e_{v}$ denoting the linear velocity error, and $e_{p}$ and $e_{\theta}$ representing the errors in position and orientation, respectively. Scaling coefficients $S_{p}$ and $S_{\theta}$, which adjust the impact of position and orientation errors, were both set to 0.5 in our experiments. Additionally, $p$ sums up to three penalties ($p_{act}$, $p_{vel}$, $p_{\omega}$) designed to discourage excessive thruster activation or reaching states with elevated linear and angular velocities. Our experimentation with various penalty configurations led us to adopt a penalty for thruster activation, Eq. (8) as well as excessive angular velocities, Eq. (9). Here, $T$ stands for an indicator function reflecting the on-off states of the thrusters.
$p_{act}$ $=0.3\sum_{i=1}^{8}T_{i}$ (8)
$p_{\omega}$ $=0.15\max(0,|\omega_{z}|-1)$ (9)
@@ III-B Simulation
Building upon our prior simulator RANS [5], we introduce enhancements to enable the platform to perform more complex tasks. RANS leverage Nvidia’s IsaacSim, specifically relying on OmniIsaacGym [25], a versatile simulator, capable of concurrently running thousands of environments. In the original RANS framework, only nominal system and environmental conditions were present. This hindered the ability of the agents to adapt to non-ideal conditions, which are usually common when using the real FP systems. To mitigate this gap, we introduce RANS v2.0 which includes the following extensions: (1) parameterized rewards and penalties, to allow easy fine-tuning of the control policies; (2) analogue kinematic model in Mujoco [26], to allow easy evaluation of both traditional and RL-based controllers in a non-Torch depended environment; (3) disturbance generation module, that allows the injection of: (a) Action Noise (AN): a random disturbance force of $\pm$ $an$ N applied to every thruster; (b) Velocity Noise (VN): $\pm$ $vn$ m/s added to the state velocities; (c) Uneven Floor (UF): $uf$ N of force, added to simulate the floor unevenness, applied to the FP body throughout the episode, either with a constant direction or through a sinusoidal generated direction; (d) Torque Disturbance (TD): $td$ Nm of torque applied to the body’s center of mass; (e) Random Thrusters Failure (RTF): a zeroing mask over the output actions to simulate one or multiple thruster failures which remains the same throughout the episode.
RANS v2.0, requires 30 minutes to train an agent on an RTX 4090. Achieving a throughput of more than 40,000 steps per second with all disturbances enabled, which is very close to its previous version. Furthermore, it enables large-scale testing by swiftly evaluating thousands of initial conditions in seconds. It offers rich visualization options, including metric tracking during training through the WandB API [27], and comprehensive evaluation metrics presented through tables and plots. The library uses the OpenAI Gym [28] format to define the RL loop, including the standard normalization of the observation space. Additionally, the integration of a ROS interface enhances the versatility of our framework, allowing easy integration and deployment of the control policies within real-world robotic systems.
@@ III-C Training Procedure
We reworked the PPO implementation from the RL Games library [29] as the foundation of our training procedure. This implementation utilizes GPU acceleration to vectorize observations and actions, enabling parallelization within the simulator by having both the simulation and the policy training residing on GPU. Our agents are designed as actor-critic networks with two hidden layers, each consisting of 128 units. This makes them light and fast enough to be ran at high frequency on embedded devices. The hyper-parameters are listed in Table IV in the appendix. The agents train in their respective environments for 2000 epochs (approximately 130M steps). For more details about the network or the PPO configuration, we invite the reader to refer to the training configuration files available along with the code release at https://github.com/elharirymatteo/RANS.
III-D Benchmark comparison with an Optimal Controller
@@
In this paper, we aim to provide a benchmark comparison between deep reinforcement learning and optimal control approaches, LQR in particular, for addressing the control problem of the floating platform in various scenarios. Our objective is not to establish the superiority of one method over the other, but rather to gain insights into the strengths and weaknesses of each approach under different environmental conditions and task requirements.
An infinite horizon discrete-time LQR controller [30] is used as a preliminary comparison with the DRL algorithm to control the FP. The LQR technique utilizes linearized dynamics to comprehensively model system behavior, providing optimal solutions with long-term stability while handling minor disturbances [31]. Their adaptability and relatively straightforward implementation have resulted in their adoption for numerous space applications [7, 32, 33]. In the case of a FP, the position, linear velocities, orientation quaternions, and angular velocities in the 2D plane are considered state variables of the system, $\mathbf{X}_{i}$. Since a FP operates at a relatively high frequency, a linearized system dynamics, defined as (10)
$\mathbf{X}_{k+1}=\mathbf{A}\mathbf{X}_{k}+\mathbf{B}\mathbf{U}_{k}$ (10)
is sufficient to predict the control output for incremental steps. The linearized system matrices, represented by $\mathbf{A}$ and $\mathbf{B}$, are the partial derivatives of the state vector at the final time step, denoted as $\mathbf{X}_{k+1}$, with respect to the current time step, $\mathbf{X}_{k}$, and the control input $\delta\mathbf{U}_{k}$, respectively. This computation leverages the central differencing technique, where the effects on the final states are evaluated in response to deliberate and minor perturbations applied to both the states and control inputs within the kinematic model simulated in Mujoco. To better account for the disturbances endured by the FP, the system matrices are updated at regular intervals. The LQR controller minimizes the cost function:
$J=\sum_{k=0}^{\infty}\mathbf{X}_{k}^{\text{T}}\mathbf{Q}\mathbf{X}_{k}+\mathbf{U}_{k}^{\text{T}}\mathbf{R}\mathbf{U}_{k}$
where $\mathbf{Q}$ and $\mathbf{R}$ are weighting matrices that penalize state errors and control outputs. Minimizing the aforementioned cost function delivers an optimal control sequence given by:
$\mathbf{U}_{k}=-\mathbf{K}\mathbf{X}_{k}$
where $\mathbf{K}$ is the control feedback gain matrix defined by:
$\mathbf{K}=(\mathbf{R}+\mathbf{B}^{\text{T}}\mathbf{P}\mathbf{B})^{-1}\mathbf{B}^{\text{T}}\mathbf{P}\mathbf{A}$
such that $\mathbf{P}$ is a positive definite matrix that is a solution for the Algebraic Riccati equation, as in:
$\mathbf{P}=\mathbf{Q}+\mathbf{A}^{\text{T}}\mathbf{P}\mathbf{A}-\mathbf{A}^{\text{T}}\mathbf{P}\mathbf{B}\mathbf{K}.$
The optimal control output, $\mathbf{U}_{k}$, is an eight-dimensional array with real numbers. Note that the control outputs correspond to the actuation of the eight thrusters on the FP, hence an alternate vector $\mathbf{\bar{U}}_{k}$ is implemented that is a least squares solution to:
$\min\;\;||\;\mathbf{\bar{U}}_{k}-\mathbf{U}^{\prime}_{k}\;||^{2}$
where $\mathbf{U}^{\prime}_{k}$ is the normalized vector of $\mathbf{U}_{k}$ with values between 0 and 1. Moreover, for $\mathbf{\bar{U}}_{k}=[u_{1},u_{2},...,u_{8}]$, each $u_{i}$ for $i\in\{1,2,...,8\}$ represents a binary variable, i.e., $u_{i}\in\{0,1\}$ signifying the actuation state of each thruster as either “on” or “off”.
III-E Laboratory Experiment Setup